OpenStudy (anonymous):

Solve the following compound inequality. Write the solution in in interval notation. -3<-2x+1over 5<= 3

OpenStudy (anonymous):

i got [-7,8) but not sure if its right

OpenStudy (anonymous):

\[-3< \frac{ -2x+1 }{ 5 }\le3\]

OpenStudy (anonymous):

that one looks better i got [-7,8) for the answer thanks for any help

OpenStudy (tkhunny):

So, \(-15 < -2x+1 \le 15\). What did you do next?

OpenStudy (anonymous):

sub 1 from both sides

OpenStudy (anonymous):

then divide both sides by -2 an change the equation signs from less than and less than or equal to...to greater than and greater than or equal to

OpenStudy (tkhunny):

\(-16 < -2x \le 14\) Divide by -2 \(8 > x \ge -7\) [-7,8) I think you have it.

OpenStudy (anonymous):

see thats what i thought, hummm my professor acted as though it was wrong but ill have to check back with her...thanks for ur help

OpenStudy (anonymous):

she mentio to keep an eye on the brackets but idk

OpenStudy (tkhunny):

Looks good to me. Go have that discussion.